Hello, I need your help!
I got a system of these three equations.
1. 0 =4a+2b+c
2. -2=9a+3b+c
3. -2= a- b+c
I don't know how to change those to find out the equation y=ax^2+bx+c.
Thanks!

Answer by stanbon(75887)

Hello, I need your help!
I got a system of these three equations.
1. 0 =4a+2b+c
2. -2=9a+3b+c
3. -2= a- b+c
I don't know how to change those to find out the equation y=ax^2+bx+c.
---------------
I used my TI-84 to get:
a = -2/3
b = 4/3
c = 0
===========================
Cheers,
Stan H.
==================